# Belo Horizonte, MG, Brazil

### v3

@@ 3.2 Bayesian Inferences
Prior distributions for all the model parameters are to be specified to complete the Bayesian model specification. For fixed effects regression coefficients ($\bm{\beta}$), we employ uncorrelated zero-mean Normal priors with marginal variances $10^{2}$. In the homoscedastic case, on $\sigma$ we place a half-t prior with $3$ degrees of freedom, denoted $t_{+}(3)$ (Gelman,, 2006). For the hereroscedastic model (2), we set independent standard normal priors for $\alpha_{0}$ and $\alpha_{1}$. The PEC dependence is encoded through $\bm{\delta}=\{\rho,\nu\}$. In geostatistics, smoothness parameters for correlation functions are typically fixed, and their estimation properties are rarely reported (Zhang,, 2004). Due to this and the scope of this paper, we specify separate models for different values of $\nu$ ensuring a positive definite PEC and select the best fit for inference. The spatial dependence parameter $\rho$ is also inherently difficult. Thus, to avoid oversmoothing, we employ an exponential prior for $\rho$ with mean $\lambda_{\rho}=-\rho_{0}/\log(p_{\rho})$, where $\rho_{0}$ is an upper-bound set to four-fifths of the largest observed Hausdorff distance, and $p_{\rho}$ is the probability that $\rho$ exceeds $\rho_{0}$. The resulting prior concentrates its mass on lower values $\rho$.
@@
$\pi(\bm{\theta},\mathbf{z}\mid\mathbf{y})\propto p(\mathbf{y}\mid\mathbf{X},\mathbf{z},\bm{\beta},\bm{\gamma})\pi(\mathbf{z}\mid\bm{\sigma},\bm{\delta})\pi(\bm{\theta}),$ (7)
@@
Parameter Estimation
The estimation of the spatial GLMM parameters is obtained from the posterior distribution defined in Equation (7). As the posterior distribution is intractable, we use the No-U-Turn (Homan and Gelman,, 2014) Markov Chain Monte Carlo (MCMC) sampler available in Stan (Team,, 2023) to draw samples from it. The No-U-Turn algorithm samples all the model parameters jointly and efficiently exploits the parameter space using automatic differentiation. Additionally, it eliminates the need for hand-tuning, making it a highly convenient sampler. We initialize the parameters with samples from their respective prior distributions. The latent random effects are initialized from a standard Gaussian distribution. The number of samples and the warm-up period of the MCMC algorithm are application dependent. We assess the convergence of the chains using the split-$\hat{R}$ diagnostic (Vehtari et al.,, 2021). Finally, the parameters’ point estimates and 95% credible intervals (CI) are obtained as, respectively, the median and percentiles ($2.5$ and $97.5$, unless otherwise stated) of the marginal MCMC samples.
Model Comparison
The assessment of goodness-of-fit (GoF) in the simulation study and applications involved two model comparison criteria: the leave-one-out information criterion (Vehtari et al.,, 2017, LOOIC), and the widely applicable information criterion (Watanabe and Opper,, 2010, WAIC). For both of them, lower values indicate a better fit. We did not use the deviance information criterion due to its assumption of a small number of parameters relative to the sample size (McElreath,, 2016, Ch. 7.4), which would require computationally intensive numerical integration of random effects in some of our models. When comparing a single candidate model to the HGP, we utilize the difference in the GoF metric $G$ between the HGP and the given model, denoted by $\Delta_{G}$. Consequently, a negative $\Delta_{G}$ suggests that the HGP model is a better fit for the data. The competing methods depend on the type of application we are working with and will be mentioned in the upcoming sections.
Predictions
Predicting spatial phenomena accurately at unobserved locations is crucial in many applications, especially in change of support settings (Gelfand et al.,, 2001). The HGP prior for spatial random effects in GLMMs offers a powerful solution. For a set of $m$ unobserved locations $\mathcal{S}^{\ast}=\{\mathbf{s}^{\ast}_{1},\ldots,\mathbf{s}^{\ast}_{m}\}$, we define an $m$-dimensional variable $\mathbf{Y}^{\ast}$ representing the response variable at these locations. Assuming covariates from Equation (4) are available, we can generate predictions accounting for uncertainty using the posterior predictive distribution of $\mathbf{Y}^{\ast}$:
$p(\mathbf{y}^{\ast}\mid\mathbf{y})=\int p(\mathbf{y}^{\ast}\mid\mathbf{z}^{\ast},\bm{\theta})p(\mathbf{z}^{\ast}\mid\mathbf{z},\bm{\theta})\pi(\bm{\theta}\mid\mathbf{y}){\rm d}\bm{\theta},$ (8)
where $\mathbf{z}^{\ast}={(z(\mathbf{s}^{\ast}_{1}),\ldots,z(\mathbf{s}^{\ast}_{m}))}^{\top}$ is a realization of the spatial random effect at the new locations, and $\bm{\theta}$ represents the model parameters. Details on sampling from this distribution are provided in Section S.1 of the Appendix Material.
@@ 4.1 Areal Data
We conducted a simulation study to assess the performance of the HGP relative to the DAGAR model, a state-of-the-art model designed explicitly for areal data (Datta et al.,, 2019), in modeling data generated from other models. We grounded our simulations in a real-world context by utilizing the 134 Intermediate Zones (IZ) north of the River Clyde within the Greater Glasgow in Scotland (Figure S.1), mirroring the application detailed in the upcoming Section 5.1. This region exhibits a range of Hausdorff distances between zones, from 0.7 to 37.7 kilometers (median: 9.4 km). Data were generated under a Poisson GLMM with a log link function, incorporating a single explanatory variable drawn from a standard normal distribution. The intercept and regression coefficients associated with this explanatory variable were fixed at 4.28 and 0.33, respectively, to mirror the parameters estimated in the real-world application (Section 5.1).
Two random effect structures were considered: the aggregated Gaussian process (AGP) and the Besag–York–Mollié (Besag et al.,, 1991, BYM), with model parameters corresponding to their fits to the real data. The AGP assumes a point-level GP aggregated at the areal unit level and is closer in structure to the HGP model. We chose AGP’s model parameters to reflect the HGP results from our real-world application (Section 5.1): a point-level marginal standard deviation ($\sigma$) of 0.19, a smoothness parameter ($\nu$) of 1, and a range parameter ($\rho$) of 2.21. The BYM model allows for spatially correlated random effects with a structure closer to that of the DAGAR model. We set its marginal standard deviation (Riebler et al.,, 2016), denoted $\sigma$, to 0.24 and the spatial dependence parameter ($\xi$) to 0.43, aligning with estimates from the real data analyses in Section 5.1. Section S.2.1 of the Appendix Material provides further details on the models used to generate data.
We generated 200 datasets from the Poisson GLMM under each of the two distinct random effect structures. For each simulated dataset, we fit two Bayesian Poisson GLMMs, the first one using an homoscedastic HGP while the second one using a DAGAR prior for the random effects. Both specified the fixed effects correctly with the single covariate. The priors for the parameters in the HGP model were set to be those presented in Section 3.2. The priors for the parameters in the DAGAR model were set to the same as those in(Datta et al.,, 2019). In particular, we set $\pi(\beta_{0})\propto 1$ and $\beta_{1}\sim N(0,1000^{2})$ for the regression parameters; a Gamma prior with shape 2 and rate 1 was placed on the conditional precision ($\phi$), and a uniform prior between 0 and 1 on the spatial dependence parameter ($\psi$). The GoF was evaluated using the criteria in Section 3.2, namely, LOOIC and WAIC, to compare which model fits the data better.
@@
Figure 3 shows the HGP consistently fitting the data better than the DAGAR model when data were generated from the AGP. The figure displays empirical cumulative distribution functions (ECDFs) of the differences ($\Delta_{G}$) in LOOIC and WAIC (left and right columns of the panel, respectively) between the models. While no formal thresholds exist for significant differences in the model comparison criteria we used, we consider differences exceeding 5 to be noteworthy, following a similar practice used with the DIC (Vranckx et al.,, 2021). For AGP-generated data (solid line), our simulation study provides strong evidence that the HGP outperforms the DAGAR in GoF. Specifically, $\Delta_{\textrm{LOOIC}}$ was below 10-10 for 80% of simulated datasets, strongly favoring the HGP. In contrast, evidence favoring DAGAR ($\Delta_{\textrm{LOOIC}}>4$) was observed in only 1% of the generated data. Interestingly, the HGP remained competitive even when we generated data from the BYM model (a structure closer to the DAGAR). While there was some evidence favoring the DAGAR, $\Delta_{\textrm{LOOIC}}$ was below 4 for more than 50% of the simulations under this scenario. These results highlight the HGP’s robustness as a disease mapping model, demonstrating superior performance when the data aligns with its structure and remains a viable option even when a specialized competitor better represents the underlying process.
In the Appendix Material (Table S.1), we also evaluated the HGP’s ability to recover true model parameters under correct specification. We assessed estimation accuracy using Mean Absolute Percentage Error (MAPE) and the frequentist coverage percentage (CP) of 95% credible intervals. All model parameters achieved CPs close to the nominal level, with the lowest being 92.5% for $\rho$, which consistently proved the most challenging to estimate. The marginal standard deviation ($\sigma$) exhibited a maximum MAPE of only 7%.
@@ 4.2 Data Fusion
Figure 4: Maps utilized in the data fusion simulation study. The left map displays the region used for fitting the models, while the right map shows the data points used to evaluate the models’ predictive accuracy.
@@
To assess the predictive performance of the HGP against standard data fusion models, we conducted a simulation study mirroring usual data fusion scenarios. Within a square region with an area of $51^{2}$, we generated a map comprising a $17\times 17$ grid-boxes of areal units (each with an area of $3^{2}$) and 35 point-referenced data points. Additionally, we created five extra spatial units for each of the following three types exclusively to evaluate predictions: larger polygons (grid boxes with area of $3^{2}$), smaller polygons (grid boxes with area of $1.5^{2}$), and point-referenced, as displayed in Figure 4. We carefully chose the study region and grid box dimensions so each polygon’s area exceeded one, aligning with typical data fusion problems. This design emulates scenarios where areal units, often derived from satellite or computer models, are more abundant than point-referenced data. Hausdorff distances between sample units in this map ranged from $0.4$ to $69.8$. We simulate the datasets from a Normal GLMM with an identity link function, no covariates, and a spatial random effect. The overall mean of this GLMM was fixed at $6.28$ to mirror the application from Section 5.2. In addition, this model introduces the small scale variance (denoted $\tau^{2}$).
@@
This section’s spatial GLMMs employ two distinct spatial random effect structures: a heteroscedastic HGP and an AGP. The HGP’s SD function $v(\mathbf{s})$ is defined as $\exp\{\alpha_{0}+\alpha_{1}\mathbbm{1}(\lvert\mathbf{s}\rvert>0)\}$, allowing for distinct variances between point-referenced and areal data. We selected HGP parameters ($\tau^{2}=0.17^{2}$, $\sigma=4.4416$, $\sigma_{a}=1.3114$, $\rho=28.45$, and $\nu=0.8$) to closely resemble the data analysis presented in Section 5.2. Note that the spatial variance for point and areal units ($\sigma$ and $\sigma_{a}$, respectively) is a one-to-one transformation of $(\alpha_{0},\alpha_{1})$$^{\top}$. We set the dependence parameter $\rho$ to 48.9. This value differs from our estimate in the real data due to a necessary rescaling based on the map used in this simulation study. For further details, refer to Section S.3 in the Appendix Material.
The AGP model assumes:
[…]
where $\{Z(\mathbf{s})\,:\,\mathbf{s}\in D\}$ is a zero-mean, second-order stationary GP with an isotropic Matérn covariance function with smoothness parameter $\nu=1$. SeeMoraga et al., (2017) for a detailed definition of this function. We simulated AGP data with parameters $\sigma^{2}=2.71^{2}$, $\rho=2.29$, and $\tau^{2}=1\times{10}^{-16}$, again mirroring the real data analysis in Section 5.2. In aggregated models, $\tau$ approaches zero as only the average of homoscedastic independent residuals is observed, which converges in probability to zero.
@@
We then generated 200 datasets from the Normal GLMM under each of the two different spatial random effects’ structures detailed next. For each simulated dataset, we fit two Bayesian Normal GLMMs, one using the aforementioned heteroscedastic HGP prior for the random effect and the other using the AGP prior. When fitting the HGP model, we used the priors highlighted in Section 3.2. In addition, we put a penalized complexity prior (Simpson et al.,, 2017) on $\tau^{-2}$, such that $\mathbbm{P}(\tau>a)=.05$, where $a$ is half the empirical SD observed in the data. Inferences and predictions were based in 2000 MCMC samples with a 1000 warm-up phase per chain with four parallel chains. The initialization of the algorithm is described in Section 3.2.
Inferences for the model induced by the spatial random effect from Equation (9) were performed using the integrated nested Laplace approximation and the stochastic partial differential equation (Lindgren et al.,, 2011, SPDE) method. The joint distribution induced by this AGP involve approximating stochastic integrals. Practitioners must define a grid with a fixed number of points within each areal unit to approximate these integrals (Wang and Furrer,, 2021). The SPDE uses a mesh to discretize the space to achieve such an approximation efficiently. The number of grid boxes per polygon will depend on the resolution of the mesh. We used two different meshes to show the inferences and, consequently, the predictions may change substantially depending on the accuracy of the approximation for the stochastic integrals in Equation (9). In the Appendix Material, Figure S.2 visualizes these meshes: a sparse mesh (left column), and a fine-resolution mesh (second column), with approximately 2 and 25 points per polygon, respectively. Henceforth, we will differentiate the AGP models by mesh density, denoting the sparse-mesh model as AGP1 and the fine-mesh model as AGP2. Their priors are detailed in Section S.4 of the Appendix Material.
@@
We assess the quality of model predictions across different scenarios using four metrics. Firstly, point and interval predictions are derived from the posterior predictive distributions (Equation (8)), using their median and the 2.5% and 97.5% percentiles, respectively. The root mean square error of prediction (RMSP) is used to evaluate point estimates. The remaining three metrics assess interval predictions: the frequentist coverage percentage of prediction intervals, denoted CPP, their width, and the interval score (Gneiting and Raftery,, 2007, IS). The latter is defined as
@@
Table 1: Out-of-sample prediction for the three models under different data generating models. “Sp. unit” indicates the geometry of the sample units for which predictions were assessed; for polygons (denoted Poly), area is additionally specified within parenthesis. Fit specifies the model fitted to the data. RMSP is the root mean squared error of prediction, CPP is the frequentist coverage percentage of the 95% prediction interval, Width is the prediction interval width, and IS is the interval score. Double headers denote the true data-generating model.
@@
Sp. unit Fit RMSP CPP Width IS RMSP CPP Width IS
Overall AGP1 1.56 67.8 3.10 17.74 1.81 57.9 2.09 31.37
AGP2 1.39 95.7 5.59 6.60 1.80 88.1 5.53 10.63
HGP 1.65 90.8 6.10 8.65 1.51 95.7 6.15 7.05
Poly ($3^{2}$) AGP1 0.78 75.8 1.91 5.79 0.89 72.5 1.94 6.96
AGP2 0.60 96.3 2.50 2.87 0.80 89.7 2.68 4.14
HGP 0.83 97.9 3.78 4.00 0.74 96.4 3.00 3.36
Poly ($1.5^{2}$) AGP1 1.33 75.6 3.56 11.73 0.90 75.6 2.12 6.47
AGP2 1.13 95.8 4.69 5.38 1.09 96.6 4.72 5.32
HGP 1.50 79.2 3.84 9.36 0.85 94.8 3.36 4.01
Points AGP1 2.56 52.2 3.85 35.70 3.64 22.8 2.20 80.70
AGP2 2.43 95.1 9.59 11.54 3.51 77.9 9.19 22.42
HGP 2.62 95.3 10.7 12.57 2.94 96.0 12.09 13.79
@@
Table 1 presents the prediction assessment metrics under both data-generating models at different types of spatial units. Let us first consider the “overall” case, that is, ignoring the types. When the AGP is the true model, the fine mesh AGP model yields the most accurate point predictions (lowest RMSP). However, the HGP model’s performance is only marginally worse, with a RMSP 18% higher. For interval predictions, the HGP model produces prediction intervals (PI) with near-nominal CPP and widths comparable to the AGP2 model, resulting in a competitive interval score (IS). Despite having the narrowest PIs on average, the sparse-mesh AGP model suffers from poor coverage and, consequently, a high IS. As expected, when the true data-generating model is the HGP, the HGP model’s predictive performance surpasses that of its competitors. The point predictions from both AGP models are remarkably similar, with the fine-mesh AGP exhibiting a RMSP only 19% higher than the correctly specified HGP. This suggests the fine-mesh AGP remains competitive in terms of point predictions, despite model misspecification. In terms of interval predictions, while the AGP2 was competitive, its sparse mesh counterpart delivered poor interval predictions. The performance of the competing method varies considerably depending on the mesh resolution, which can significantly impact prediction accuracy. In contrast, our model eliminates the need for arbitrary mesh selection and consistently delivers highly competitive predictions, enhancing prediction accuracy.
Table 1 further reveals a consistent decline in prediction accuracy for all models as the area of the spatial units decreased. For AGP-generated data, the AGP2 model’s RMSP for point-referenced data was approximately four times higher than for the polygons with bigger area (i.e., $3^{2}$). Similarly, the HGP model’s RMSP increased by a factor of 3.15 for the same comparison (point-referenced vs largest polygon). For both the HGP and AGP2, CPP remained near the nominal level for all the spatial units. However, HGP’s CPP dropped below nominal level for smaller polygons (area of $1.5^{2}$). This is likely due to variance underestimation: while the AGP-generated data’s variance is inversely proportional to spatial units, the fitted HGP model only distinguish between points and polygons, not their area. Regarding the IS, both our model and the AGP2 model showed IS values roughly three times higher for point-referenced data compared to the largest polygons. This aligns with the simulation parameters. The AGP1 model’s IS performance deteriorated substantially when decreasing the area of the spatial units.
Under the HGP data-generating process, the HGP model consistently outperformed competitors on all assessment metrics except PI width. Although point predictions were broadly similar across models, the HGP provided remarkably superior interval predictions for all spatial unit types, mantaining CPP near the nominal level (see Table 1). In contrast, the AGP1 model exhibited a CPP of only 22.8% for point-referenced data. The poor coverage for both AGP models with point-referenced data is further reflected in their IS, which were 5.9 (AGP1) and 1.6 (AGP2) times higher than the HGP. On the other hand, the AGP2 model remained competitive with the HGP for polygons interval predictions, while the sparse AGP performed poorly.
Table S.3 in the Appendix Material presents the results for parameter estimation under the correct specification of the the HGP model. We used the same metrics as in Section 4.1 to assess the parameters estimation. The MAPE for $\rho$ was around $10.9\%$ while for $\sigma_{a}$ it was $9.21\%$. Similarly to what is observed with geostatistical models, the small scale standard deviation parameter, namely $\tau$, was the most challenging parameter to be estimated. Unlike the areal simulation, the coverage of some of the variance parameters was consistently below the nominal level of $95\%$. This results are expected, as the HGP inherits the weak identifiability of these parameters from geostatistical models under a normal likelihood (Zhang,, 2004). Predictions remain trustworthy despite this issue.

### v4

@@ 3.2 Bayesian Inferences
Prior distributions for all the model parameters are to be specified to complete the Bayesian model specification. For fixed effects regression coefficients ($\bm{\beta}$), we employ uncorrelated zero-mean Normal priors with marginal variances $\mathrm{v}_{\beta}$. In the homoscedastic case, on $\sigma$ we place a half-t prior with $3$ degrees of freedom, denoted $t_{+}(3)$ (gelman2006prior). For the hereroscedastic model (2), we set independent Normal priors for $\alpha_{0}$ and $\alpha_{1}$ with marginal variances $\mathrm{v}_{\alpha}$. Hereafter, we set the marginal variances as $\mathrm{v}_{\beta}=10^{2}$ and $\mathrm{v}_{\alpha}=1$ unless otherwise stated. The PEC dependence is encoded through $\bm{\delta}=\{\rho,\nu\}$. In geostatistics, smoothness parameters for correlation functions are typically fixed (zhang2004inconsistent; datta2016hierarchical). Due to this and the scope of this paper, we specify separate models for different values of $\nu$, selected based on the empirical assessment to ensure positive definiteness of the induced correlation matrices, and select the best fit for inference. The spatial dependence parameter $\rho$ is also inherently difficult to estimate. Therefore, to avoid oversmoothing, we set an exponential prior in the same spirit of the penalized complexity (PC) prior (simpson2017penalising; fuglstad2019constructing) for $\rho$. Specifically, we employ an exponential prior for $\rho$ with mean $\lambda_{\rho}=-\rho_{0}/\log(p_{\rho})$, where $\rho_{0}$ is an upper-bound set to four-fifths of the largest observed Hausdorff distance, and $p_{\rho}$ is the probability that $\rho$ exceeds $\rho_{0}$. The resulting prior concentrates its mass on lower values $\rho$. Note that, the resulting prior is not exactly a PC prior since it is not placed on the Kullback-Leibler divergence with a simpler model.
@@
$\pi(\bm{\theta},\mathbf{z}\mid\mathbf{y})\propto\pi(\mathbf{y}\mid\mathbf{X},\mathbf{z},\bm{\beta},\bm{\gamma})\pi(\mathbf{z}\mid\bm{\sigma},\bm{\delta})\pi(\bm{\theta}),$ (4)
@@
The estimation of the spatial GLMM parameters are obtained from the posterior distribution defined in Equation (4). Since this distribution is intractable, we draw samples from it using the No-U-Turn Markov chain Monte Carlo (MCMC) sampler in Stan (hoffman2014no; stan2023rstan). Unless otherwise stated, we initialize the parameters with samples from their respective prior distributions. The latent random effects are initialized from a standard Gaussian distribution. The number of samples and the warm-up period of the MCMC algorithm are application dependent. We assess the convergence of the chains using the split-$\hat{R}$ diagnostic (vehtari2021rank). Finally, the parameters’ point estimates and 95% credible intervals (CI) are obtained as, respectively, the median and percentiles ($2.5$ and $97.5$, unless otherwise stated) of the marginal MCMC samples.
We assess goodness-of-fit (GoF) using the leave-one-out (vehtari2017practical, LOOIC) and widely applicable (watanabe2010asymptotic, WAIC) information criteria, where lower values indicate a better fit. When comparing a single candidate model to the HGP, we utilize the difference in the GoF metric $G$ between the HGP and the given model, denoted by $\Delta_{G}$. Consequently, a negative $\Delta_{G}$ suggests that the HGP model is a better fit for the data. The competing methods depend on the type of application we are working with and will be mentioned in the upcoming sections.
Predicting spatial phenomena accurately at unobserved locations is crucial in many applications. The HGP prior for spatial random effects in GLMMs offers a powerful solution. For a set of $m$ unobserved locations $\mathcal{S}^{\ast}=\{\mathbf{s}^{\ast}_{1},\ldots,\mathbf{s}^{\ast}_{m}\}$, we define an $m$-dimensional variable $\mathbf{Y}^{\ast}$ representing the response variable at these locations. Assuming covariates from Equation (3) are available, we can generate predictions accounting for uncertainty using the posterior predictive distribution of $\mathbf{Y}^{\ast}$:
$\pi(\mathbf{y}^{\ast}\mid\mathbf{y})=\int\pi(\mathbf{y}^{\ast}\mid\mathbf{z}^{\ast},\bm{\theta})\pi(\mathbf{z}^{\ast}\mid\mathbf{z},\bm{\theta})\pi(\bm{\theta}\mid\mathbf{y})\,{\rm d}\mathbf{z}^{\ast}\,{\rm d}\mathbf{z}\,{\rm d}\bm{\theta},$ (5)
where $\mathbf{z}^{\ast}={(z(\mathbf{s}^{\ast}_{1}),\ldots,z(\mathbf{s}^{\ast}_{m}))}^{\top}$ is a realization of the spatial random effect at the new locations, and $\bm{\theta}$ represents the model parameters. Details on sampling from this distribution are provided in Section S.3 of the Supplementary Material.
@@ 4.1 Areal Data
We conducted a simulation study to assess the performance of the HGP relative to the DAGAR model, a state-of-the-art model designed explicitly for areal data (datta2019spatial), in modeling data generated from other models. For details about the DAGAR model, we refer to the Appendix. We grounded our simulations in a real-world context by utilizing the 134 Intermediate Zones (IZ) north of the River Clyde within the Greater Glasgow in Scotland (Figure S.5), mirroring the application detailed in the upcoming Section 5.1. This region exhibits a range of Hausdorff distances between zones, from $0.7$ to $37.7$ kilometers (median: $9.4$ km). Data were generated under a Poisson GLMM with a log link function, incorporating a single explanatory variable drawn from a standard normal distribution. The intercept and regression coefficients associated with this explanatory variable were fixed at $4.28$ and $0.33$, respectively, to mirror the parameters estimated in the real-world application (Section 5.1).
Two random effect structures were considered: the Aggregated Gaussian Process (AGP) and the Besag–York–Mollié (besag1991bayesian, BYM), with model parameters corresponding to their fits to the real data. Details about those models and their parameters are given in the Appendix, Section A. The AGP assumes a point-level GP aggregated at the areal unit level and is closer in structure to the HGP model. We chose AGP’s model parameters to reflect the HGP results from our real-world application (Section 5.1): a point-level marginal standard deviation ($\sigma$) of $0.19$, a smoothness parameter ($\nu$) of $1$, and a range parameter ($\rho$) of $2.21$. The BYM model allows for spatially correlated random effects with a structure closer to that of the DAGAR model. We set its marginal standard deviation (riebler2016intuitive), denoted $\sigma$, to $0.24$ and the spatial dependence parameter ($\xi$) to $0.43$, aligning with estimates from the real data analyses in Section 5.1. In Section S.5.2 from the Supplementary Material, we detail how to simulate data from the AGP.
We generated 200 datasets from the Poisson GLMM under each of the two distinct random effect structures. For each simulated dataset, we fit two Bayesian Poisson GLMMs, the first one using an homoscedastic HGP while the second one employing a DAGAR prior for the random effects. Both specified the fixed effects correctly with the single covariate. The priors for the parameters in the HGP model were set to be those presented in Section 3.2. The priors for the parameters in the DAGAR model were set to the same as those in (datta2019spatial); For more details, see the Supplementary Material, Section S.4.1. The GoF was evaluated using the criteria in Section 3.2. In the Supplementary Material, Section S.4.2, we also evaluated the HGP’s ability to recover true model parameters under correct specification.
@@
Figure 2 shows the HGP consistently fitting the data better than the DAGAR model when data were generated from the AGP. The figure displays empirical cumulative distribution functions (ECDFs) of the differences ($\Delta_{G}$) in LOOIC and WAIC (left and right columns of the panel, respectively) between the models. While no formal thresholds exist for significant differences in the model comparison criteria we used, we consider differences exceeding 5 to be noteworthy, following a similar practice used with the DIC (vranckx2021stability). For AGP-generated data (solid line), our simulation study provides strong evidence that the HGP outperforms the DAGAR in GoF. Specifically, $\Delta_{\textrm{LOOIC}}$ was below 10-10 for 80% of simulated datasets, strongly favoring the HGP. In contrast, evidence favoring DAGAR ($\Delta_{\textrm{LOOIC}}>4$) was observed in only 1% of the generated data. Interestingly, the HGP remained competitive when we generated data from the BYM model (a structure closer to the DAGAR). While there was some evidence favoring the DAGAR, $\Delta_{\textrm{LOOIC}}$ was below 4 for more than 50% of the simulations under this scenario. These results highlight the HGP’s robustness as a disease mapping model, demonstrating superior performance when the data aligns with its structure and remains a viable option even when a specialized competitor better represents the underlying process.
@@ 4.2 Data Fusion
To assess the predictive performance of the HGP against standard data fusion models, we conducted a simulation study mirroring usual data fusion scenarios. Within a square region with an area of $51^{2}$, we generated a map comprising a $17\times 17$ grid-boxes of areal units (each with an area of $3^{2}$) and 35 point-referenced data points. Additionally, we created five extra spatial units for each of the following three types exclusively to evaluate predictions: larger polygons (grid boxes with area of $3^{2}$), smaller polygons (grid boxes with area of $1.5^{2}$), and point-referenced, as displayed in Figure S.6 from the Supplementary Material. We carefully chose the study region and grid box dimensions so each polygon’s area exceeded one. This design emulates scenarios where areal units, often derived from satellite or computer models, are more abundant than point-referenced data. Moreover, the grid-boxes of different sizes used to assess the predictions have the purpose of investigating how the predictions deteriorate by decreasing the size of the spatial units. Hausdorff distances between sample units in this map ranged from $0.4$ to $69.8$. We simulate the datasets from a Normal GLMM with an identity link function, no covariates, and a spatial random effect. The overall mean of this GLMM was fixed at $6.28$ to mirror the application from Section 5.2. In addition, this model introduces the small scale variance (denoted $\tau^{2}$).
This section’s spatial GLMMs employ two distinct spatial random effect structures: a heteroscedastic HGP and an AGP. The HGP’s SD function $v(\mathbf{s})$ is defined as $\exp\{\alpha_{0}+\alpha_{1}\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)\}$, allowing for distinct variances between point-referenced and areal data. While the AGP is a zero-mean second-order stationary GP with an isotropic Matérn covariance function with smoothness parameter $\kappa=1$. See the Appendix for further details on the AGP model, and moraga2017geostatistical for a detailed definition of the Matérn covariance function.
@@
We selected HGP parameters ($\tau^{2}=0.17^{2}$, $\sigma=4.4416$, $\sigma_{a}=1.3114$, $\rho=28.45$, and $\nu=0.8$) to mirror what was estimated in the data analysis presented in Section 5.2. Note that the spatial variance for point and areal units ($\sigma$ and $\sigma_{a}$, respectively) is a one-to-one transformation of $(\alpha_{0},\alpha_{1})$$^{\top}$. We set the dependence parameter $\rho$ to 48.9. This value differs from our estimate in the real data due to a necessary rescaling based on the map used in this simulation study. For further details, refer to the Supplementary Material, Section S.5.1. We simulated AGP data with parameters $\sigma^{2}=2.71^{2}$, $\rho=2.29$, and $\tau^{2}=1\times{10}^{-16}$, again mirroring the real data analysis in Section 5.2. In aggregated models, $\tau$ approaches zero as only the average of homoscedastic independent residuals is observed, which converges in probability to zero.
@@
We then generated 200 datasets from the Normal GLMM under each of the two different spatial random effects’ structures detailed above. For each simulated dataset, we fit three Bayesian Normal GLMMs for each simulated dataset, one using the aforementioned heteroscedastic HGP prior for the random effect and other two using the AGP prior. When fitting the HGP model, we used the priors highlighted in Section 3.2. In addition, we put a penalized complexity prior (simpson2017penalising) on $\tau^{-2}$, such that $\mathbbm{P}(\tau>a)=.05$, where $a$ is half the empirical SD observed in the data. Inferences and predictions were based in 2000 MCMC samples with a 1000 warm-up phase per chain with four parallel chains. Inferences for GLMMs with an AGP model are detailed in Section S.5.3 from the Supplementary Material.
An important aspect of the AGP model lies in the approximation of the stochastic integrals’ from Equation (6). They are usually approximated numerically (fuentes2005model). In order to show the inferences and the predictions may change substantially depending on the accuracy of these approximations, we considered a two different grids to compute those integrals numerically: a sparse- (denoted, AGP1) and a fine-resolution (denoted, AGP2) grid, with approximately 2 and 25 points per polygon, respectively. Their priors are detailed in Section S.5.3 of the Supplementary Material. In Supplementary Material, Section S.5.4, we also present the results for parameter estimation under the correct specification of the the HGP model.
@@
We assess the quality of model predictions across different scenarios using three metrics. Firstly, point and interval predictions are derived from the posterior predictive distributions (Equation (5)), using their median and the 2.5% and 97.5% percentiles, respectively. The root mean square error of prediction (RMSP) is used to evaluate point estimates. The remaining two metrics assess interval predictions: the frequentist coverage percentage of prediction intervals, denoted CPP, and the interval score (gneiting2007strictly, IS). The latter is defined as
@@
Table 1: Out-of-sample prediction for the three models under different data generating models. Fit specifies the model fitted to the data. RMSP is the root mean squared error of prediction, CPP is the frequentist coverage percentage of the 95% prediction interval, and IS is the interval score. Double headers denote the true data-generating model.
@@
Fit RMSP CPP IS RMSP CPP IS
AGP1 1.56 67.8 17.74 1.81 57.9 31.37
AGP2 1.39 95.7 6.60 1.80 88.1 10.63
HGP 1.65 90.8 8.65 1.51 95.7 7.05
@@
Table 1 presents the prediction assessment metrics under both data-generating models. When the AGP is the true model, the fine mesh AGP model yields the most accurate point predictions (lowest RMSP). However, the HGP model’s performance is only marginally worse, with a RMSP 18% higher. For interval predictions, the HGP model produces prediction intervals (PI) with near-nominal CPP and competitive IS. The AGP1 model suffers from poor coverage and, consequently, a high IS. As expected, when the true data-generating model is the HGP, the HGP model’s predictive performance surpasses that of its competitors. The point predictions from both AGP models are remarkably similar, with the fine-mesh AGP exhibiting a RMSP only 19% higher than the correctly specified HGP. This suggests the fine-mesh AGP remains competitive in terms of point predictions, despite model misspecification. In terms of interval predictions, while the AGP2 was competitive, its sparse mesh counterpart delivered poor interval predictions. The performance of the competing method varies considerably depending on the mesh resolution, which can significantly impact prediction accuracy. In contrast, our model eliminates the need for arbitrary mesh selection and consistently delivers highly competitive predictions, enhancing prediction accuracy. Additional analyses assessing predictions across different spatial resolutions are presented in the Supplementary Material, Section S.5.4.